A LABORATORY AND THEORETICAL STUDY OF PROTONATED CARBON DISULFIDE, HSCS^{+}

\underline{M. C. MCCARTHY,} P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138; JEREMIAH J. WILKE AND HENRY F. SCHAEFER III, Center for Computational Chemistry, University of Georgia, 1004 Cedar St, Athens, GA 30602.

The rotational spectrum of protonated carbon disulfide, HSCS^{+}, has been detected in the centimeter-wave band in a molecular beam by Fourier transform microwave spectroscopy. Rotational and centrifugal distortion constants have been determined from transitions in the K_{a} = 0 ladder of the normal isotopic species, HS^{13}CS^{+}, and DSCS^{+}. The present assignment agrees well with high-level coupled cluster calculations of the HSCS^{+} structure, which, like earlier work, predict this isomer to be the ground state on the HCS_{2}^{+} potential energy surface; HCSS^{+}, an isomer with C_{2v} symmetry, is predicted to lie more than 20 kcal/mol higher in energy. Other properties of HSCS^{+} including its dipole moment, vibrational frequencies, and infrared intensities have also been calculated at the CCSD(T)/cc-pwCVQZ level of theory. Because carbon disulfide possesses a fairly large proton affinity, and because this nonpolar molecule may plausible exist in astronomical sources, HSCS^{+} is a good candidate for detection with radio telescopes in the sub-millimeter band where the stronger b-type transitions of this protonated cation are predicted to lie.